OpenStudy (breathlessreaper):

I need help badly, please help? 3. Using your function f(x) from question 2. Demonstrate and explain how to find the average rate of change between year 3 and 5, and between year 5 and 7. Explain what the average rate of change represents to the frog population. (x) = 200(1.03^4 was my equation I used before. How would I figure this out??? I don't understand.

OpenStudy (anonymous):

http://openstudy.com/study#/updates/5292f6d3e4b00bed1904e17f Try that

OpenStudy (phi):

you should "tweak" your function f(x)= 200(1.03)^x <--- notice you want an "x" in the equation. 200(2.03)^4 is not a function of x because there is no x!

OpenStudy (anonymous):

A reaper can help XD

OpenStudy (breathlessreaper):

@phi OHH okay, that makes more sense...

OpenStudy (breathlessreaper):

@AkaylaM Thank you

OpenStudy (phi):

find the average rate of change between year 3 and 5 that means find (f(5) - f(3))/(5-3) f(5) is short for : in your equation, replace x with 5, and do the arithmetic For example, using f(x)= 200(1.03)^x , we would do f(5)= 200(1.03)^5 if you type into google 200(1.03)^5 = it will give you the answer 231.854815 do the same for f(3) then subtract the two numbers. that is how much your function changed. divide by 5-3 (i.e. divide by 2) to get the average rate of change per year.